OpenStudy (anonymous):

Determine a series form for the power series:

OpenStudy (anonymous):

-(sqrt(2)/3)x^-3 + (sqrt(6)/5)x^-5 - (sqrt(12)/7)x^-7 + (sqrt(20)/9)x^-9 - .............

OpenStudy (anonymous):

Clarify: Im asking to write those terms in one form. ex. series n=1 t infin of x^n.

OpenStudy (anonymous):

Shall I post the origional quest? might be easier

OpenStudy (experimentx):

yeah ... you can post

OpenStudy (anonymous):

\[\int\limits_{0}^{.4}\sqrt{1+X ^{4}} dx\] |Error|<5x10^6

OpenStudy (experimentx):

is |Error| = n ??

OpenStudy (anonymous):

no.

OpenStudy (anonymous):

find a series for that integral, then use Rn<integral of f(x)

OpenStudy (anonymous):

Im gussing that is series representation ?

OpenStudy (experimentx):

\[ 1+\sum_{n=1}^{\infty} \frac{(-1)^{n+1}x^{4n+1}}{4n(4n+1)} \] Integration of http://www.wolframalpha.com/input/?i=expand+sqrt%281%2Bx^4%29